A small body is projected up a rough(coefficent of friction=0.5)inclined plane,(inclined angle=60 degree)with a speed 10 metre per second. How far along the plane it moves up before coming to rest?? g=10 ms^-2

Grade:

1 Answers

Kunjal sheth
8 years ago
We know that work done by all forces = change in K.E.----1
And work done = net force x displacement 
Therefore
W.D. = (umgcos@ +mgsin@)x displacement
By 1 
(umgcos@ +mgsin@)x displacement (s)=(1/2)mvv
(umgcos(60) + mgsin(60)) x s=(1/2)m(100)
((1/2)(10)(1/2)+(10)(√3/2)) x s =50
S= (20)/2√3+1